Status Report

ISS On-Orbit Status 21 Feb 2002

By SpaceRef Editor
February 21, 2002
Filed under , ,

All ISS systems continue to function nominally, except as noted previously or below.

The crew received emphatic words of appreciation and thanks for their “picture perfect” EVA yesterday (official duration: 5:47:23), which “made everyone here on the ground very happy”.

During post-EVA activities last night, when back in the cabin, the crew noted a strong foul smell, emanating from the Airlock (A/L). It was traced to the regeneration process of Metox canisters and assessed by the ground as harmless, possibly stemming from mildew on the canisters which had been left in the oven from 7A, i.e., since July 2001. The regen process was terminated and the A/L hatch closed. The intermodular ventilation (IMV) between A/L and the rest of the station was deactivated, and the Lab TCCS (trace contaminant control system) has been scrubbing the Lab and Node atmosphere all day today. The crew spent much of the day in the Russian segment (RS), and all activities that would have required them to be in the US segment (USOS) were cancelled. PMA-1 hatch to Node is partially closed. Since Dan and Carl had much off-duty time today, this impacted only a few tasks, such as post-EVA deconfiguring of the A/L (including Metox regeneration) and station repress from Progress. [Recyclable Metox (metal oxide) canisters have replaced the old expendable LiOH (lithium hydroxide) canisters as carbon dioxide (CO2) removal system in the EMU/spacesuits. During use, CO2 is absorbed by them and later removed through a special valve opening by “baking” (heating), which takes place in a special oven in the A/L.]

Before breakfast, FE-1 Walz and FE-2 Bursch were scheduled for the Russian post-EVA MedOps assessments MO-8 (body mass measurement) and MO-9 (biochemical urine test).

Later, the two Flight Engineers completed the important post-EVA PuFF (pulmonary function in flight) experiment session, as well as U.S. post-EVA health status checks, assisting each other as “medical officer”.

Early in the morning, during Daily Orbits (DO) 1 & 2, Progress M1-7 (6P), at the SM aft end, performed a small reboost of the station. For this purpose, ISS attitude was placed under RS control at midnight and then changed moderately by RS thrusters. The two short maneuver burns of about 3 minutes each (delta-V: 2.8 m/sec total) served as test of the eight thrusters to be used for the upcoming larger ISS reboost. The burns were under MCC-M/TsUP control via pre-stored (time-tagged) onboard commands; they raised mean station altitude by 4.7 km. Since they were generally perigee burns, most of the altitude increase happened at apogee, which went up by 7.7 km (perigee: 1.6 km). This increased the eccentricity (ellipticity) of the orbit.

ISS was then pointed to the attitude required for the planned BINS (Russian strap-down inertial guidance system) test corrections, which CDR Onufrienko completed during orbital night using first the Puma sighting device (DO3), and 90 minutes later the VP-2 pilot sight (on DO4). His reference stars, taken from the EKZ 5.1 star map laptop program, were Alpha Canis Majoris (Sirius), Beta Canis Majoris (Mirzam), Gamma Velorum and Zeta Puppis.

The planned repressurization of the ISS cabin air to 760 mmHg with fresh oxygen (O2) from Progress 6Pâs high-pressure O2 tankage system (SRPK) was postponed due to the “smell event”. This repress procedure conserves electric power and service life of the Russian Elektron O2 generator. [Progress M1 uses N2O4 (nitrogen tetroxide] as oxidizer in its propulsion system, but also carries about 40 kg of O2 as resupply cargo.]

Onufrienko also completed a number of routine tasks, including continued checkout of the faulty IK0501 gas analyzer in the SM for its CO2 readings, air sample collection using the Russian AK-1M sampler (for Freon content analysis), and monitoring the microbial contamination of the ISS atmosphere with the “Ecosphere” MedOps experiment SZM-MO-21 set up on 2/12.

Later in the day, Yuri removed the old amateur radio packet module at the ISS Ham station in the FGB and replaced it with a new packet module, followed by a test. He configured it for its checkout by the ground operator at W5RRR in Houston and stowed the old unit for future use. The new transceiver allows ground amateur operators to use it as a beacon or message exchange system to world-wide ground amateurs.

A ham radio pass was scheduled at 10:00 am EST with a German technical high school for surveyors, named “Peter Anich”, located in Bolzano, Northern Italy.

After Bursch and Walz had configured the DDCU (dc-to-dc converter unit) umbilicals during yesterday’s EVA for checkout by the ground, it was discovered that they were not in the proper configuration. The spacewalkers were therefore asked to leave the cables connected and CIDs (circuit interrupt devices) closed, to enable further work. This will be done by software uploads, currently being planned. It is expected that the DDCUs will be operating properly in ample time for 8A.

A caution alert was annunciated early yesterday morning for one of the BCDUs (battery charge/discharge units) of the P6 2B power channel, caused by a bit flip in the BCDU EEPROM (electrically erasable reprogrammable read-only memory) chip. The BCDU was successfully refreshed from the ground after the EVA and is operating nominally.

The transition to the new software release for mission/stage 8A will begin tonight during crew sleep, starting with CDH (command and data handling) MDM data dumps to MCC-H, followed by CDH 8A software transition/step-up continuing all day tomorrow, along with reboots, reconnections and ACS (attitude control system) handovers. Ku-band will be unavailable until about 10:30 am EST. S-band pointing will also not be available from the GNC MDM (guidance, navigation and control computer) but will be conducted in “manual” pointing mode (using time-tag commands from MCC-H) for proper antenna alignment with the TDRS satellite in sight. A detailed contingency plan has been developed in the event of problems with communications during the transition. By tomorrow evening (7:00 pm EST), MCC-H expects to have two C&C (command and control) MDMs and one GNC MDM running the 8A flight software. S-band pointing will then be available from the GNC (primary only) and able to operate in Auto tracking again.

CEO (crew earth observation) targets today were E. Mediterranean Dust and Smog (high pressure is beginning to break down in this region, but as ISS crossed the Egyptian coast, crew was to look to the right of track toward the Sinai, Israel and Lebanon for oblique views of aerosol accumulation there), Tigris-Euphrates, Turkey (an excellent pass over the new reservoirs built or under construction to manage increasingly critical water resources. Crew was to map these lakes [250 and 350 mm lens]; recent snowfall should enhance their contrast and features of the terrain), European Smog (as ISS crossed the heel of Italy, crew was to use 50 and 110 mm lenses for oblique [left] shots of smog plumes over the Adriatic Sea), Gulf of St. Lawrence (visibility was deteriorating from the west on this pass, but crew may still have been able to document pack ice formation near Cape Breton Island and the southwestern coast of Newfoundland), Gulf of St. Lawrence (of interest on this second pass over this area: mapping ice patterns along the NW coast of the island of Newfoundland), and Lakes of the Eastern Sierra Watershed (crew to use this brief break in the winter storm series to document the Sierra snow pack and possibly Lake Tahoe and Pyramid Lake the NE).

U.S. and Russian Segment Status (as of 2:08 pm EST today):

Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):

  • Elektron O2 generator is On (16 Amp mode, the lowest possible setting).
  • Vozdukh CO2 scrubber is in MANUAL cycle mode #5 (vacuum pump failed). U.S. CDRA CO2 scrubber is in Standby.
  • BMP Harmful Impurities unit: Both absorbent beds (Filters #1 & #2) in Purify mode.
  • SM Working Compartment: Pressure (mmHg) — 752, temperature (deg C) — 26.3, ppO2 (mmHg) — 160.9, ppCO2 (mmHg) — 2.5 (? see Notes below and above)

  • SM Transfer Compartment: Pressure (mmHg) — 752, temperature (deg C) 21.0; ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.

  • FGB Cabin: Pressure (mmHg) — 756, temperature (deg C) — 23.0; ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.

  • Node: Pressure (mmHg) — 743.25, temperature (deg C) — 23.2 (shell); ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.

  • U.S. Lab: Pressure (mmHg) — 745.62, temperature (deg C) — 22.5, ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a;

  • Joint Airlock (Equip. Lock): Pressure (mmHg) — 746.12, temperature (deg C) — 22.0; shell heater temp (deg C) — 22.8, ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.

  • PMA-1: Shell heater temp (deg C) — 23.7.

  • PMA-2: Shell heater temp (deg C) — 22.8.

(Note: Partial pressures ppO2 and ppCO2 in U.S. segment (USOS) not available because MCA [major constituent analyzer] is failed and in Extended Life mode [= a state that preserves mass spectrometer vacuum but produces no pp data]. Russian GA (gas analyzer) readings of ppCO2 in the SM are invalid).

Electrical Power Systems (EPS):

  • Beta Gimbal Assembly (BGA) 2B in Autotrack mode, BGA 4B in Autotrack mode (during XPOP).

  • SM batteries: all batteries (8) in “partial charge” mode.

  • FGB battery #5 is offline; battery #4 is cycling; all other (4) batteries are in “partial charge” mode.

  • Plasma Contactor Unit PCU-1 in Standby mode; PCU-2 in Standby mode.

Thermal Control Systems:

  • Air conditioner SKV-1 is Off (Freon leak). SKV-2 is On.

Command & Data Handling Systems:

  • C&C-2 MDM is prime, C&C-3 is back-up, and C&C-1 is in standby.

  • GNC-2 MDM is prime; GNC-1 is back-up.

  • LA-1, LA-2 and LA-3 MDMs are all operating.

  • APS-1 (automated payload switch #1) and APS-2 are both On.

  • SM Terminal Computer (TVM): 3 redundant lanes (of 3) operational.

  • SM Central Computer (TsVM): 2 redundant lanes (of 3) operational.

Communications Systems:

  • S-band is operating nominally.

  • Ku-band is operating nominally in open loop pointing mode.

  • Audio subsystem operating nominally.

  • Video subsystem operating nominally.

  • MCOR (medium-rate communications outage recorder) operating nominally.


  • SSRMS/Canadarm2 at Port Stow position, with Keep Alive power on both strings.

  • RWS (robotics workstations) are Off.

ISS Orbit (as of this morning, 9:36am EST):

  • Mean altitude — 385.8 km

  • Apogee — 392.2 km

  • Perigee — 379.5 km

  • Period — 92.3 min.

  • Inclination (to Equator) — 51.63 deg

  • Eccentricity — 0.0009403

  • Orbits per 24-hr. day — 15.61

  • Altitude increase — 4.7 km (mean) in last 24 hours

  • Solar Beta Angle: -51.3 deg (magnitude increasing)

  • Revolutions since FGB/Zarya launch (Nov. â98) — 18600

  • Current Flight Attitude — XPOP (x-axis perpendicular to orbit plane [yaw: ~0 deg, pitch: 5 deg., roll: 0 deg])

For more on ISS orbit and naked-eye visibility dates/times, see

SpaceRef staff editor.